Shift report 20000627 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Jun 27 2000 - 23:46:25 EDT

  • Next message: http account: "Shift report 20000628 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Jun 27 23:46:24 2000, Shift: 16:00-24:00
    Supervisor: FR, with: ZIS
    Log book#1 pages 90 - 90
    
    Shift Summary:
    ------------------------------------------------------------
     - Machine was down at the beginning of the shift (power dip).
       At 18:30 MCR started machine setup, then they tried to
       ramp to g=70 but the ramp failed. At 22:30, another
       unsuccessful ramp (yellow beam did not survive).
       Current status is "Ramping magnets down".
    
     - Runs 1563 and 1564 are pedestal runs.
    
     - No detector problems to report. TPC anode voltages are OFF.      
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Jun 27 2000 - 23:50:18 EDT